# Cost-Optimal Microservices Deployment with Cluster Autoscaling and Spot Pricing

### v1

@@ Abstract
Microservices architecture has been established as an ideal software architecture for cloud-based software development and deployment, offering many benefits such as agility and efficiency. Microservices are often associated with containers and container orchestration systems for deployment, as containerization provides convenient tools and techniques for resource management, including the automation of orchestration processes. Among the factors that make the cloud suitable for commercial software deployment, transient pricing options like AWS Spot Pricing are particularly attractive as they allow consumers to significantly reduce cloud costs. However, the dynamic nature of resource demand and the abrupt termination of spot VMs make transient pricing challenging. Nonetheless, containerization and container orchestration systems open new avenues to optimize the cost of microservices deployments by leveraging spot pricing on the public cloud while achieving application and business goals.
We propose SpotKube, an open-source, Kubernetes-based, application-aware, genetic algorithm-based solution for cost optimization, which autoscales clusters for microservices-based applications hosted on public clouds with spot pricing options. SpotKube analyzes application characteristics and recommends the optimal configuration for resource allocation to the cluster. It consists of an elastic cluster autoscaler powered by an optimization algorithm that ensures cost-effective microservices deployment while meeting application performance requirements and handling abrupt termination of nodes, thereby minimizing the impact on system availability. We implement and evaluate SpotKube with representative microservices-based applications in a real public cloud setup, demonstrating the effectiveness of our approach against alternative optimization strategies.
@@
Microservice, Cost Optimization, Transient Pricing, Cluster AutoScaling
@@ I INTRODUCTION
Microservices architecture has gained immense popularity in recent years due to its ability to provide agility, flexibility, and efficiency to software development and deployment in the cloud environment [1]. This architectural pattern enables the creation of complex applications by breaking them down into smaller, more manageable services that can be developed, deployed, and scaled independently.
Container orchestration is an essential component when running containerized microservice workloads, as it automates much of the operational effort required to run those workloads and services. However, a cost-optimized container orchestration solution is necessary to minimize the overall financial costs for the consumer. By using transient pricing models, such as AWS Spot Pricing, which offers a dynamic pricing model, consumers can further minimize operational costs.
Optimal resource utilization is a critical factor for optimizing the costs associated with deploying microservices applications. Due to the dynamic and complex nature of these applications and the abrupt termination nature of spot VMs, achieving this optimization presents a considerable challenge. Consequently, a thorough characterization of the application, coupled with a comprehensive analysis of resource usage, is vital.
Numerous studies have focused on solving distinct use cases such as microservice application characterization, cost optimization through transient pricing models for microservice applications, and cost-optimized deployment strategies in public cloud environments. However, there exists untapped potential to develop a comprehensive framework that facilitates the deployment of microservice workloads in a public cloud environment, incorporating application characteristic analysis to deliver maximum monetary benefits to consumers via cost-optimized cluster autoscaling and transient pricing while simultaneously ensuring robust performance and gracefully handling the abrupt termination of spot instances.
As a result, we propose SpotKube, an open-source Kubernetes-managed service that optimizes the deployment cost of microservices applications. It achieves this through an application characterization tool that determines the optimal deployment configuration to meet user-defined SLOs and cost-optimized cluster autoscaling. SpotKube utilizes nodes from public cloud environments, leveraging transient pricing models such as AWS Spot Pricing to further minimize operational costs.
In this paper, we implement and evaluate SpotKube with representative microservices-based applications on a real public cloud setup. We compare the effectiveness of our approach against alternative optimization strategies and demonstrate that SpotKube significantly reduces cloud costs while ensuring that application performance requirements are met.
Key Contributions of the paper:
@@
    Measurement driven reactive approach for cost-optimal microservices application deployment using spot instances
@@
    An Elastic Cluster Autoscaler, powered by an optimization engine based on genetic algorithm, aims to find the optimal balance between cost and performance.
@@
The remainder of this paper is organized as follows: Section II offers an overview of related work on microservices architecture and cost optimization in public cloud environments. Section III details the SpotKube architecture, including a comprehensive explanation of its implementation. Sections IV and V present the experimental setup, followed by the results and their analysis. Section VI discusses and analyzes the key findings of SpotKube, while Section VII outlines future work and concludes the paper.
@@ II RELATED WORKS
This section presents a comprehensive literature review on different aspects of deploying microservices-based applications in a public cloud environment in a cost-optimized manner, including application characterization, cost modeling, auto-scaling, and container scheduling.
@@ II-A Application Characterization
Application characterization is a crucial step in deploying microservices-based applications in cloud environments. The process involves analyzing the performance of an application and identifying its optimal configurations to ensure optimal resource utilization and application performance.
Jindal et al. [2] introduced a novel tool, Terminus, which employs a methodology encompassing series of load tests to assess resource usage and calculate Microservice Capacity (MSC), defined as the maximum rate of HTTP or gRPC requests that can be served while adhering to Service Level Objectives (SLOs). One of the main drawbacks of this approach is that it can be time-consuming to conduct a series of load testing to estimate MSC.
Han et al. [3] proposed a framework for microservices placement in cloud-native computing environments that utilizes empirical profiling to identify and respond to workload characteristics. The framework involves conducting profiling experiments on selected workloads to derive delicate resource requirements and then performing microservices placement using a greedy-based heuristic algorithm that considers application performance based on the resource requirements. The proposed framework is practical and utilizes empirical profiling to identify and respond to workload characteristics, which has the potential to improve the efficiency and performance of microservices placement.
Overall, the two papers emphasize the importance of application characterization and its impact on microservices-based application deployment in cloud environments.
@@ II-B Cost modeling
A. Baldominos Gómez et al. [5] proposes the use of various machine learning techniques to predict the prices of Elastic Cloud Compute (EC2) spot instances offered by Amazon Web Services. The study uses two historical datasets and data extracted from the EC2 API to train separate regression models based on the instance type. The results show that forecasting performance varies significantly between instance types, with some models achieving almost perfect predictions while others barely outperform the baseline prediction of average price. The study concludes by introducing a Prediction-as-a-Service system, which is deployed in the cloud. This study only analyzes the last three months of data and the performance of the models may vary over time, yet the model has performed well for certain instance types.
@@ II-C Auto-scaling and scheduling containers
Kubernetes is an open-source system for orchestrating and managing the lifecycle of containerized applications [6]. [7, 8, 9] propose extended components for Kubernetes to support auto-scaling and cost-optimized scheduling of containers.
Sami et al. [7] proposed an agent named FScaler, which utilizes reinforcement learning to horizontally scale the container’s instances and schedule the placement of newly created instances based on defined cost functions after studying the change in resource availability.
Rajkumar Buyya et al. [8] developed a custom scheduler, rescheduler and autoscaler as extended components for Kubernetes. The scheduler places incoming tasks on a cluster resource with at least the amount of CPU and memory resources requested by the task in an optimized manner. Then the autoscaler scale out based on the cluster’s workload to meet the resource demand and scale in to utilize the underutilized VMs to reduce the cost. The rescheduler consolidates applications into fewer VMs when possible, avoiding costly scaling-out operations or encouraging scale-in operations if the tasks can not be rescheduled.
M. Yan et al. [9] developed a tool to improve the horizontal auto-scaling policy of Kubernetes using SARSA based multi-agent parallel training module. This module predicts the task arrival rate and guarantees the assurance of SLA (Service Level Agreement). The complexity and resource usage of the task scheduling algorithm is comparatively high, which is an overhead to the overall performance.
@@
Shridhar G. Domannal et al. [11] propose an efficient cost-optimized scheduling algorithm for a Bag of Tasks (BoT) using the Modified Particle Swarm Optimization technique. The authors present a Neural Network Based backpropagation algorithm to predict spot instances values to reduce the cost further.
Gu, Haihua, et al. [12] aim to schedule a suite of microservice tasks while minimizing the total renting cost of virtual machines from the cloud platform for a microservice-based workflow application that includes both the on-demand and the spot instances cost. The authors proposed a dynamic task scheduling heuristic algorithm (DSMT) to determine the number of VMs, the container configurations on those VMs, and the tasks allocated to the containers. Spot Instance Monitor is responsible for reallocating the tasks on the out-of-bid failure VMs to on-demand instances by considering both the rental cost and the sub-deadlines of interrupted tasks.
Bento, Araujo, Barbose, et al. [13] hss propose a method called Cost-Availability Aware Scaling (CAAS) which achieves cost savings primarily through a novel approach to scaling pods within a microservices architecture, utilizing a multi-objective optimization strategy that balances both cost and availability. They intergrated the Amdahl’s law with the Generalized Logistic Function (GLF), to model the non-linear variations in system availability in response to changes in workload and the number of replicas. By optimizing the number of pod replicas based on cost and availability, CAAS efficiently manages resource allocation, which leads to significant cost reductions compared to traditional autoscaling methods based on CPU utilization, like the Kubernetes Horizontal Pod Autoscaler (HPA).
@@ III-A Overview
The underlying motivation of the proposed system is to identify the most cost-efficient and application-specific allocation of resources to a Kubernetes cluster while ensuring compliance with user-defined Service Level Objectives (SLOs) and preserving optimal application performance. Provisioning the public cloud infrastructure, characterizing the application, determining the most cost-effective allocation of resources, and dynamically scaling cluster resources based on the application requirements are significant challenges encountered by this approach. In response to these challenges, we present an architecture, as depicted in Fig 1, that addresses them effectively.
@@
In summary, the user is required to provide three essential inputs, namely user configurations that incorporate SLOs, provisioning configurations, and a microservices application. Subsequently, the user must conduct load testing using an application characterization tool, and the results are then utilized by an analytical model to calculate the optimal number of replicas that can be deployed in the cluster while adhering to the user-defined SLOs. This replica count is subsequently utilized by the optimization engine, which determines the most cost-effective resource allocation to the cluster via the node allocator. The Helm service then deploys the application based on the configuration provided by the optimization engine. Furthermore, an elastic scaler has been implemented to facilitate the dynamic scaling of cluster resources while handling the abrupt termination of spot instances.
Detailed explanations of each challenge, solution, and implementation follow.
III-B Provisioning the Public Cloud Infrastructure
One of the important tasks in the SpotKube pipeline is provisioning the public cloud infrastructures based on the user-provided configuration. In order to facilitate this process, the Provisioner has been designed to automate the provisioning workflow.
Provisioner: Provisioning is a critical component of deploying microservice-based applications on public cloud, as it involves building the necessary environment in public cloud environments for deploying the microservice application and running SpotKube services. Provisioner automates this process, ensuring that the user’s configurations are implemented consistently across all public cloud environments. The provisioner is responsible for creating the necessary infrastructure, including Virtual Private Clouds (VPCs), security groups, network interfaces, and management nodes.
III-C Application Modeling
The process of application modeling represents a significant challenge due to the dynamic nature of each microservices application. In addition, it is important to note that the approach adopted by SpotKube is limited to CPU-intensive applications. In order to analyze the characteristics of the application and determine initial required number of pods (replicas) that need to be deployed, SpotKube has consist of two primary components which are Application Characterization and Analytical Model. There are certain assumptions and constraints that we used during modeling the application
@@ III-C Application Modeling
Application Characterization: The application characterization tool comprised multiple steps. Firstly, the microservice-based application needs to be subjected to load testing using Locust[14], a popular open-source tool that provides programmatically accessible APIs for simulating realistic user behavior and traffic patterns. Following this, CPU metrics are collected using Prometheus, and time series data are retrieved from Grafana API to capture the application’s performance characteristics.
To automate the load-testing workflow, SpotKube has provided a script that takes care of the entire process. Initially, the user deploys the microservices application using Minikube and the provided configuration, followed by running the aforementioned script. The duration of the load test is determined by the application, with the test running until the failure rate is surpass a specific threshold. Finally, the tool outputs the time series data which includes the load test results, providing insight into the application’s performance characteristics.
It is important to mention that deploying pods with the same resource limits on different instance types results in almost the same performance [2]. Hence regardless of the underneath infrastructure, load test results can be used for further analysis.
Analytical Model: By leveraging metrics collected through the application characterization tool and the user-provided configurations, the analytical model analyzes the behavior of the application. The model then determines the required number of pods that need to be deployed to the cluster, based on the Service Level Objectives (SLOs) specified by the user. Note that this pod count is the minimal number of pods that need to be initially deployed without violating the user defined SLOs.
The analytical model is composed of two key components, namely the Metric Analyzer and the Pod Calculator. The Metric Analyzer takes the metrics collected by the load test, which include CPU usage per load and failure rates. The primary objective of the Metric Analyzer is to determine the maximum load(request per seconds) that a pod can handle without any performance degradation.
@@
Figure 2 present data on CPU utilization and the corresponding failure rate obtained through load testing a prime number calculation application. Metrics related to CPU usage were collected using Prometheus[15]. As illustrated in Figure 2, there is a notable decline in CPU usage beyond a certain point as illustrated in the vertical line, coinciding with an increase in failure rate. This inflection point marks the application’s threshold for maximum load capacity, beyond which it experiences crashes.
@@ III-D Cost optimal Resource Allocation
Monetary cost optimization of deploying microservices applications is one of the vital research areas that we address in this paper. Due to many reasons such as the complexity, dynamicity, elasticity, and resource utilization requirements of microservice architecture, cost optimization while maintaining user-defined performance has become a more demanding task.
@@
Predictor: The predictor is utilized to predict the spot price for each predefined node type unless the user-defined them. By analyzing the last three months’ time series data, which is fetched in real-time from the AWS spot pricing history API [16], the predictor will train and predict the spot price in real-time. Since there is a very high variability in the prices depending on type [5], rather than using a pre-trained model, Prophet [17] is used to forecast the spot price in a more real-time manner.
Optimizer: The optimizer leverages the predicted spot prices to identify the best node combination that corresponds to the given workload, utilizing a brute force, a greedy approach or a genetic algorithm based Pareto optimization. The brute force approach will walk through the whole search space which is computationally expensive while the greedy approach tries to find out the optimal node configuration by using heuristics like recommending the most cost-effective node by considering the current workload. Genetic Algorithm based Pareto optimization tries to find the Pareto front of the two conflicting objectives, cost and the performance. Then this node configuration will be fed into the node allocator to allocate nodes to the cluster.
Genetic algorithm-based Pareto optimization solutions have demonstrated the best trade-off between cost and performance, which is one of the vital research outcomes of our study. The following diagram 3 illustrates the flow of the optimization approach based on NSGA-II optimization.[18].
@@
The fitness function plays a key role in identifying the optimal set of nodes that minimize total costs while maximizing the number of nodes. By maximizing the number of nodes and allocating mix of nodes belongs to different availability zones and instance families, the system’s availability is maintained even if there are sudden node terminations. This topic will be discussed in more detail in the discussion section.
In the fitness function, we ensure node sufficiency by comparing the resource availability of each node combination with the resource requirements of the currently running pods or workload.
Cost Model: The cost model calculates the cost of a selected set of nodes and returns the cost per unit time (1 hour) to the given pod count. This function definition, is inspired by and adapted from foundational concepts discussed in the related works section (II.B).
The total cost $C(m)$ for utilizing cloud services during a time period $m$ is given by the following equation:
@@
$C_{m,i}^{Ser}$ is calculated using the real-time spot price predictor model that we implemented for each spot instance type selected by the genetic algorithm-based optimizer.
Node Allocator: Based on the configuration provided by the optimization engine the node allocator requests spot instances from the spot market and assign them to the private subnet in the virtual private cloud inside the public cloud which is provisioned by the provisioner. Once the spot instances are acquired, the node allocator proceeds to configure the master node and other worker nodes, these worker nodes are then assigned to the cluster where the microservices application is deployed.
III-E Cluster Auto Scaling
Microservices applications are highly dynamic and require efficient management and automation to ensure optimal performance. Container orchestration tools like Kubernetes have emerged as effective solutions for handling such applications. Additionally, cluster auto-scaling has become an essential component of container orchestration to maintain availability and scalability based on the application’s demands.
Cost optimization in cluster auto-scaling is crucial for consumers to reduce the deployment costs. Maintaining system availability, especially with the abrupt termination of spot instances, is a challenging task for the elastic scaler. SpotKube’s elastic scaler addresses these challenges, ensuring cost-efficient auto-scaling while maintaining application availability.
Elastic Scaler: Elastic Scaler operates by periodically notifying the optimization engine about the current configuration, enabling the cluster to adjust to the application’s needs. The elastic scaler comprises two primary components, namely the Worker API and the Scaler Engine. The lightweight Worker API continuously communicates with the Kubernetes API server and compares any changes in the current configuration with the initial configuration. Upon detecting a valid change, it will invoke the Scaler Engine. The Scaler Engine will collect more data from the API server and apply certain heuristics to determine if the configuration change is significant. If the change is deemed significant, the optimization engine will be invoked with the current configuration which includes the current pod count.

### v2

# SpotKube: Cost-Optimal Microservices Deployment with Cluster Autoscaling and Spot Pricing

@@ Abstract
Microservices architecture, known for its agility and efficiency, is an ideal framework for cloud-based software development and deployment, particularly when integrated with containerization and orchestration systems that streamline resource management. However, the cost of cloud computing remains a critical concern for many organizations, prompting the need for effective strategies to minimize expenses without compromising performance. While cloud platforms like AWS offer transient pricing options, such as Spot Pricing, to reduce operational costs, the unpredictable demand and abrupt termination of spot VMs introduce considerable challenges. By leveraging containerization alongside intelligent orchestration, it is possible to optimize microservices deployment costs while maintaining performance requirements.
We present SpotKube, an open-source, Kubernetes-based solution that employs a genetic algorithm for cost optimization. Designed to dynamically scale clusters for microservice applications on public clouds using spot pricing, SpotKube analyzes application characteristics to recommend optimal resource allocations, ensuring deployments remain cost-effective without sacrificing performance. Its elastic cluster autoscaler adapts to changing demands, gracefully managing node terminations to minimize disruptions in system availability. Evaluations conducted using real-world public cloud setups demonstrate SpotKube’s superior performance and cost efficiency compared to alternative optimization strategies.
@@
Microservices, Cost Optimization, Genetic Algorithm, Transient Pricing, Cluster Auto Scaling
@@ I INTRODUCTION
Microservices architecture has gained widespread popularity for its agility, flexibility, and efficiency in cloud-based software development [1]. By breaking complex applications into manageable services, this approach allows for independent development, deployment, and scaling.
However, the operational cost of running containerized microservices can be significant, and a cost-effective container orchestration solution, combined with transient pricing models like AWS Spot Pricing is essential.
Moreover, optimal resource utilization is critical to optimizing costs in microservices deployment. Yet, the dynamic nature of these applications, coupled with the unpredictable termination of spot instances, makes achieving this optimization challenging. Therefore, a thorough characterization of applications and careful resource analysis are essential.
Numerous studies have focused on solving distinct use cases such as microservice application characterization [2], cost optimization through transient pricing models for microservice applications [3], and cost-optimized deployment strategies in public cloud environments [4]. However, there remains untapped potential for a comprehensive framework that facilitates the deployment of microservice workloads in a public cloud environment, incorporating application characteristic analysis to deliver maximum monetary benefits to consumers via cost-optimized cluster autoscaling and transient pricing while simultaneously ensuring robust performance.
We propose SpotKube, an open-source Kubernetes-managed service that optimizes the deployment cost of microservices applications. SpotKube optimizes deployment costs through an application characterization tool that determines optimal configurations based on user-defined SLOs and genetic algorithm-based cost-optimized cluster autoscaling, leveraging transient pricing models like AWS Spot Pricing in public cloud environments.
In this paper, we evaluate SpotKube with representative microservices applications in a public cloud setup and compare its effectiveness with other optimization strategies. Our results show that SpotKube significantly reduces costs while meeting performance requirements.
Key Contributions:
@@
    Measurement-driven reactive approach for cost-optimal microservices application deployment using spot instances
@@
    An Elastic Cluster Autoscaler, powered by an optimization engine based on a genetic algorithm, aims to find the optimal balance between cost and performance.
@@
The paper is organized as follows: Section II reviews related work. Section III details the SpotKube architecture and implementation. Sections IV describe the experimental setup and results. Section V discusses key findings, and Section VI concludes the paper and outlines future work.
@@ II RELATED WORKS
This section presents a comprehensive literature review on the following aspects including microservice application characterization, cost modeling, auto-scaling, and container scheduling.
@@ II-A Application Characterization
Application characterization involves analyzing the performance of an application and identifying its optimal configurations to ensure optimal resource utilization and application performance.
Jindal et al. [2] introduced a novel tool, Terminus, which employs a methodology encompassing a series of load tests to assess resource usage and calculate Microservice Capacity (MSC), defined as the maximum rate of HTTP or gRPC requests that can be served while adhering to Service Level Objectives (SLOs).
Han et al. [5] proposed a framework for microservices placement in cloud-native computing environments that utilizes empirical profiling to identify and respond to workload characteristics. The framework involves conducting profiling experiments on selected workloads to derive resource requirements and then performing microservices placement using a greedy-based heuristic algorithm.
@@ II-B Cost modeling
A. Baldominos Gómez et al. [7] employed machine learning techniques to predict AWS EC2 spot instance prices using historical datasets and EC2 API data. The results indicated significant variability in forecasting performance across different instance types.
@@ II-C Auto-scaling and scheduling containers
Kubernetes is an open-source system for orchestrating and managing the lifecycle of containerized applications [8] and recent studies have proposed extended components for Kubernetes to support auto-scaling and cost-optimized scheduling of containers. [9, 10, 11]
Sami et al. [9] proposed an agent named FScaler, which utilizes reinforcement learning to scale the container’s instances horizontally and schedule the placement of newly created instances based on defined cost functions after studying the change in resource availability.
Rajkumar Buyya et al. [10] developed custom extensions for Kubernetes, including a scheduler, rescheduler, and autoscaler. The scheduler optimally places tasks on cluster resources with the required CPU and memory. The autoscaler adjusts the number of VMs based on workload, scaling out to meet demand and scaling in to reduce costs by utilizing underutilized VMs
M. Yan et al. [11] developed a tool to improve the horizontal auto-scaling policy of Kubernetes using the SARSA-based multi-agent parallel training module. This module predicts the task arrival rate and guarantees the assurance of SLA (Service Level Agreement). The complexity and resource usage of the task scheduling algorithm is comparatively high, which is an overhead to the overall performance.
@@
Shridhar G. Domannal et al. [13] propose an efficient cost-optimized scheduling algorithm for a Bag of Tasks (BoT) using the Modified Particle Swarm Optimization technique. The authors present a Neural network-based backpropagation algorithm to predict spot instance values to reduce the cost further.
Gu, Haihua, et al. [3] propose DSMT, a dynamic task scheduling algorithm to minimize VM renting costs for microservice workflow applications. A critical component, the Spot Instance Monitor, handles abrupt terminations by reallocating tasks from failed spot instances to on-demand instances, balancing rental costs and sub-deadlines of interrupted tasks.
Bento, Araujo, Barbose, et al. [4] have proposed CAAS a method that optimizes cost savings while maintaining availability in a microservices architecture. CAAS uses a multi-objective optimization strategy to balance cost and availability by scaling pods. CAAS achieves significant cost reductions compared to traditional pod autoscaling methods like Kubernetes Horizontal Pod Autoscaler (HPA)
@@ III-A Overview
The primary objective of the proposed system is to determine the most cost-efficient and application-specific allocation of resources to a Kubernetes cluster while ensuring adherence to user-defined Service Level Objectives (SLOs) and maintaining optimal application performance.
The architecture, as depicted in Figure 1, summarizes the overall flow of our proposed approach and it can be divided into the following key sections.
@@
Pre Deployment
  • Microservice Application Characterization: This phase involves analyzing the microservice application through a load test and collecting application-specific metrics, such as CPU usage and memory usage per given Requests Per Second (RPS).
  • Calculation of Initial Resource Requirements: This phase entails mapping the application-specific CPU and memory requirements to the user-defined SLOs and calculating the initial number of pods required for deployment. This is done by the Analytical Model.
  • Cost optimal Resource Allocation: This phase involves mapping the required pod count to available spot instances, calculating the cost-effective spot instance combination, and attaching it to the cluster. This process will be managed by the Optimization Engine.
Currently, the cluster is prepared for initial deployment. Following the successful deployment of the application, scaling the cluster becomes essential to ensure consistent performance and efficient cost management. The next phase of our approach is specifically designed to address these requirements.
Post Deployment
  • Dynamic Scaling of the Cluster: This phase is iterative and it involves dynamically attaching or detaching cluster resources based on incoming traffic, in a cost-effective manner while handling the abrupt termination of spot instances, using the Elastic Scaler and Optimization Engine.
The following sections provide detailed explanations of each step, including the assumptions made, the solution, and the implementation process.
@@
III-B Microservice Application Characterization
The process of application characterization represents a significant challenge due to the dynamic nature of each microservices application. In addition, it is essential to note that the approach adopted by SpotKube is limited to CPU-intensive applications.
The application characterization tool comprised multiple steps. Firstly, the microservice-based application needs to be subjected to load testing using Locust[14]. Following this, CPU metrics are collected using Prometheus [15].
To automate the load-testing workflow, SpotKube has provided a script that takes care of the entire process. Initially, the user deploys the microservices application using Minikube and the provided configuration, followed by running the aforementioned script. The duration of the load test is determined by the application, with the test running until the failure rate surpasses a specific threshold. Finally, the tool outputs the load test results which includes CPU usage per given RPS.
It is important to mention that deploying pods with the same resource limits on different instance types results in almost the same performance [2]. Hence regardless of the underlying infrastructure, load test results can be used for further analysis.
III-C Calculation of Initial Resource Requirements
By leveraging metrics collected through the application characterization tool the analytical model calculates the application-specific resource requirement and then determines the required number of pods that need to be deployed to the cluster, based on the Service Level Objectives (SLOs) specified by the user. Note that this pod count is the minimal number of pods that need to be initially deployed without violating the user-defined SLOs.
@@ III-C Calculation of Initial Resource Requirements
The analytical model is composed of two key components, namely the Metric Analyzer and the Pod Calculator. The Metric Analyzer takes the metrics collected by the load test, which include CPU usage per load and failure rates. The primary objective of the Metric Analyzer is to determine the maximum load(request per second) that a pod can handle without any performance degradation.
Figure 2 presents data on CPU utilization and the corresponding failure rate obtained through load testing of a prime number calculation application. Note that we have only deployed a single pod for this testing. As illustrated in Figure 2, there is a notable decline in CPU usage beyond a certain point as illustrated in the vertical line, coinciding with an increase in failure rate. This inflection point marks the application’s threshold for maximum load capacity that a pod can handle, beyond which it experiences crashes.
@@ III-D Cost optimal Resource Allocation
This paper addresses the critical challenge of optimizing monetary costs while maintaining user-defined performance for microservices applications.
@@
Predictor: The predictor forecasts the spot price for each predefined node type unless specifically defined by the user. It analyzes time series data from the past three months, retrieved in real-time from the AWS Spot Pricing History API [16], and continuously updates its model to predict spot prices. Given the high variability across instance types [7], Prophet [17] is employed for its flexibility in modeling complex seasonality and its ability to adjust dynamically to real-time data. This method outperforms static, pre-trained models, offering more accurate short-term forecasts, especially in environments with rapid price fluctuations.
Optimizer: The optimizer leverages the predicted spot prices to identify the best node combination that corresponds to the given workload, utilizing a brute force, a greedy approach, or a genetic algorithm-based Pareto optimization.
NSGA-II[18] based pareto optimization solutions have demonstrated the best trade-off between cost and performance, which is one of the vital research outcomes of our study. The following diagram 3 illustrates the flow of the optimization approach.
@@
The fitness function identifies the optimal set of nodes that minimize costs and maximize the number of nodes, ensuring system availability through diverse node allocation across availability zones and instance families. It ensures node sufficiency by comparing each node combination’s resource availability with the current workload’s requirements. This topic will be further discussed in the discussion section.
Cost Model: The cost model calculates the cost of a selected set of nodes and returns the cost per unit time (1 hour) to the given pod count.
@@
$C_{m,i}^{Ser}$ is calculated using the real-time spot price predictor model that we implemented.
According to the SpotKube approach, currently, the cluster is ready for deployment. Once the application is deployed, based on the incoming traffic cluster needs to be dynamically scaled which is explained in detail in the next section.
III-E Dynamic Scaling of the Cluster
Microservices applications require efficient management and automation for optimal performance, and container orchestration tools like Kubernetes are essential for handling them. SpotKube’s elastic scaler addresses the challenges of cost optimization in cluster auto-scaling and maintaining system availability, particularly with the abrupt termination of spot instances.
Elastic Scaler: Elastic Scaler operates by periodically notifying the optimization engine about the current configuration, enabling the cluster to adjust to the application’s needs. The elastic scaler continuously communicates with the Kubernetes API server and compares any changes in the current configuration with the initial configuration. Upon detecting a valid change it will collect more data from the API server and apply certain heuristics to determine if the configuration change is significant. If the change is deemed significant, the optimization engine will be invoked with the current configuration.